# Effects of tidal charge on Blandford – Znajek process around braneworld black holes

### v1

@@ I Introduction
In general, one can determine the BZ power for a given rotating BH via a regular solution of the force-free Maxwell equations in the BH spacetime. By means of perturbative techniques, the extraction rate was initially established to exhibit a quadratic scaling with the BH angular velocity, $\Omega_{\text{H}}$. However, both analytical Camilloni:2022kmx and numerical Tchekhovskoy:2009ba studies suggest that the full BZ power expression requires the inclusion of a high-spin factor, typically of the form $f(\Omega_{\text{H}})=1+a\Omega_{\text{H}}^{2}+b\Omega_{\text{H}}^{4}+\cdots$. For a rapidly rotating BH, the higher-order correction terms become relevant. So far, the formula of $f(\Omega_{\text{H}})$ for a monopolar magnetosphere in Kerr spacetime has been analytically extended to the sixth order at most Camilloni:2022kmx , while the authors in Ref. Tchekhovskoy:2009ba note the fact that high-spin factor is actually independent of the magnetospheric geometry. Recently, Camilloni et al. Camilloni:2023wyn investigated the BZ jets in the Kerr-MOG spacetime, a rotating BH solution in scalar-tensor-vector gravity, and explicitly derived its $f(\Omega_{\text{H}})$ up to orders $\mathcal{O}(\Omega_{\text{H}}^{2})$, which reveals that those expansion coefficients ($a,b,\dots$) are theory-dependent. To be specific, the high-spin factor of a Kerr-MOG BH is always smaller than that of a Kerr BH with the same mass and angular velocity. More importantly, breaking the parameter degeneracy in $\Omega_{\text{H}}$ enables the function $f(\Omega_{\text{H}})$ to serve as a new probe for testing possible deviations from general relativity (GR) in the strong-field regime Camilloni:2023wyn . This motivates us to explore the BZ power in other non-Kerr cases.
@@ II Rotating braneworld black hole
where $M$, $a$ and $b$ are BH mass, specific angular momentum and tidal charge, respectively. It should be noticed that the parameter $b$, which has the dimension of length squared, may be either positive or negative, unlike its counterpart $Q^{2}$ in Kerr-Newman metric. As expected, the solution (II) admits the Killing vectors
@@
Similarly, a negative value of $b$ extends the radius of ergosphere, implying that braneworld BHs with negative tidal charges are supposed to be more energetic objects Aliev:2005bi . We emphasize, however, that for energy extraction a larger $r_{\text{e}}$ does not inherently yield a greater amount or higher efficiency, as both of which in certain scenarios depend on the size of the ergoregion. For instance, it has been shown in deOliveira:2020lzp that the difference between $r_{\text{e}}$ and $r_{+}$ decreases with the increase of b-b, thus resulting in a smaller superradiant amplification factor for massless scalar waves in non-extremal cases.
@@ III.1 Force-free electrodynamics
In addition, the field lines must smoothly across the critical surfaces defined by $\eta^{\mu}\eta_{\mu}=0$, namely
@@ III.2 Static background solutions
It is difficult to directly solve Maxwell’s equations in a rotating background. However, for small spin $a$, an analytic magnetosphere may be obtained by means of BZ perturbative techniques Blandford:1977ds . To this end, we first need a leading-order solution as a seed.
@@ IV Blandford – Znajek rates
To the best of our knowledge, Eqs. (72) and (75) are also valid for other neutral, single-parameter modified Kerr BHs with a monopolar magnetosphere. Moreover, in the absence of the high-spin factor, the remaining $\dot{E}=2\pi\Omega_{\text{H}}^{2}/3$ and $\dot{L}=4\pi\Omega_{\text{H}}/3$ are confirmed to be exactly the same as the leading order results obtained in standard GR Camilloni:2022kmx . That is to say, the BZ mechanism cannot distinguish a Kerr BH from a braneworld one under consideration unless the high-spin factor (IV) is measured, due to the degeneracy of the spin parameter $a$ and the tidal charge $b$ in Eq. (9). Indeed, the degeneracy in the leading-order term of the BZ power has been confirmed in a general stationary and axisymmetric BH spacetime Konoplya:2021qll .
What is really of concern, as suggested in Ref. Camilloni:2023wyn , is the fractional deviation of $f_{\beta}$ with respect to its GR limit $f_{0}$, or equivalently, is the relative difference in BZ power between a Kerr and a braneworld BH with the same mass $M$ and angular velocity $\Omega_{\text{H}}$,
@@
It is clear from Fig. 6 that a positive $\beta$ always leads to a negative $\Delta\dot{E}$, and therefore tends to reduce the relative energy extraction rate. With a high-spin factor accurate to the second order in $\Omega_{\text{H}}$, the effect of positive tidal charges can suppress the BZ power of a braneworld BH by up to almost $15.2\%$ compared to that of a Kerr BH. Conversely, a negative tidal charge acts to boost the power by $66.5\%$ at most from a purely theoretical perspective. Such an enhancement, however, can only be achieved at $\beta\approx-58.8$, a value that is likely not allowed in realistic astrophysical contexts. For example, observations of the BH shadow constrain the dimensionless tidal charge to $\beta=-0.1^{+0.6}_{-0.5}$ Banerjee:2019nnj ; Banerjee:2022jog ; Krishnendu:2024jkj , corresponding to a maximum $\Delta\dot{E}$ of $11.8\%$.
Figure 6: Percentage-scaled $\Delta\dot{E}$ as a function of the dimensionless angular velocity $\omega_{\text{H}}\equiv M\Omega_{\text{H}}$ for different values of $\beta$, where the independent variable is bounded by $\sqrt{1-\beta}/(2-\beta)$ according to Eq. (9). Since there are no allowed configurations outside the dashed gray line, $\Delta\dot{E}$ reaches its minimum of approximately $15.2\%$ at $\omega_{\text{H}}=0.392$ and $\beta=0.765$. The inset shows the envelope at extremely negative tidal charges, revealing a maximum enhancement $\Delta\dot{E}_{\text{max}}\approx 66.5\%$ for $\omega_{\text{H}}=0.127$ with $\beta=-58.818$.
@@
which behaves qualitatively similarly to Eq. (76), as presented in Fig. 7.
Figure 7: Relative difference in angular momentum extraction rate, $\Delta\dot{L}$, between Kerr and braneworld Kerr BHs with identical $M$ and $\Omega_{\text{H}}$. The function (77) exhibits a maximum of $33.7\%$ at $\omega_{\text{H}}=0.143$, $\beta=-45.602$, and a minimum of $8.3\%$ at $\omega_{\text{H}}=0.401$, $\beta=0.749$.
@@ V Conclusions
For Kerr and braneworld BHs sharing the same mass $M$ and angular velocity $\Omega_{\text{H}}$, it is found that the BZ power is suppressed by the positive tidal charge, with a maximum reduction of approximately $15.2\%$, whereas a negative tidal charge can enhance the power by $66.5\%$ at most theoretically. A similar trend also holds for the angular momentum extraction rate, which is suppressed by a positive tidal charge and enhanced by a negative one, with the extremal values of $8.3\%$ and $33.7\%$, respectively. These results are helpful in placing independent constraints on the tidal charge parameter and the extra dimension size through future horizon-scale observations.

### v2

@@ I Introduction
In general, one can determine the BZ power for a given rotating BH via a regular solution of the force-free Maxwell equations in the BH spacetime. By means of perturbative techniques, the extraction rate was initially established to exhibit a quadratic scaling with the BH angular velocity, $\Omega_{\text{H}}$. However, both analytical Camilloni:2022kmx and numerical Tchekhovskoy:2009ba studies suggest that the full BZ power expression requires the inclusion of a high-spin factor, typically of the form $f(\Omega_{\text{H}})=1+a\Omega_{\text{H}}^{2}+b\Omega_{\text{H}}^{4}+\cdots$. For a rapidly rotating BH, the higher-order correction terms become relevant. So far, the formula of $f(\Omega_{\text{H}})$ for a monopolar magnetosphere in Kerr spacetime has been extended to the sixth order at most Camilloni:2022kmx, and the resulting analytic BZ power expression shows good agreement with both general-relativistic particle-in-cell and magnetohydrodynamic simulations over the full spin range Meringolo:2025bdu. It is also noteworthy that the high-spin factor is actually independent of the magnetospheric geometry Tchekhovskoy:2009ba. Recently, Camilloni et al. Camilloni:2023wyn investigated the BZ jets in the Kerr-MOG spacetime, a rotating BH solution in scalar-tensor-vector gravity, and explicitly derived its $f(\Omega_{\text{H}})$ up to orders $\mathcal{O}(\Omega_{\text{H}}^{2})$, which reveals that those expansion coefficients ($a,b,\dots$) are theory-dependent. To be specific, the high-spin factor of a Kerr-MOG BH is always smaller than that of a Kerr BH with the same mass and angular velocity. More importantly, breaking the parameter degeneracy in $\Omega_{\text{H}}$ enables the function $f(\Omega_{\text{H}})$ to serve as a new probe for testing possible deviations from general relativity (GR) in the strong-field regime Camilloni:2023wyn. This motivates us to explore the BZ power in other non-Kerr cases.
@@ II Rotating braneworld black hole
where $M$, $a$ and $b$ are BH mass, specific angular momentum and tidal charge, respectively. It should be noticed that the parameter $b$, which has the dimension of length squared, may be either positive or negative, unlike its counterpart $Q^{2}$ in Kerr-Newman metric. Recently, additional low-energy solutions to the Randall-Sundrum II model have been numerically derived by virtue of the $AdS_{5}/CFT_{4}$ dual theory Figueras:2011gd; Biggs:2021iqw. Notably, the rotating BHs Biggs:2021iqw exhibit four-dimensional behavior and approach the Kerr metric on the brane in the limit of large BHs, while tends toward a five-dimensional Myers-Perry BH with a single non-zero rotation parameter aligned with the brane in the limit of the small ones. Therefore, the solution (II) in the case of the tidal charge $b=0$ is consistent with the limit of large BHs in Biggs:2021iqw.
As expected, the solution (II) admits the Killing vectors
@@
Similarly, a negative value of $b$ extends the radius of ergosphere, implying that braneworld BHs with negative tidal charges are supposed to be more energetic objects Aliev:2005bi.
@@ III.1 Force-free electrodynamics
In addition, the field lines must smoothly cross the critical surfaces defined by $\eta^{\mu}\eta_{\mu}=0$, namely
@@ III.2 Static background solutions
It is difficult to directly solve Maxwell’s equations in a general rotating background. However, for small spin $a$, an analytic magnetosphere may be obtained by means of BZ perturbative techniques Blandford:1977ds. To this end, we first need a leading-order solution as a seed.
@@ IV Blandford – Znajek rates
To the best of our knowledge, Eqs. (72) and (75) are also valid for other neutral, single-parameter modified Kerr BHs with a monopolar magnetosphere. Moreover, in the absence of the high-spin factor, the remaining $\dot{E}=2\pi\Omega_{\text{H}}^{2}/3$ and $\dot{L}=4\pi\Omega_{\text{H}}/3$ are exactly the same as the leading-order results obtained in standard GR Camilloni:2022kmx. That is to say, the BZ mechanism cannot distinguish a Kerr BH from a braneworld one under consideration unless the high-spin factor (IV) is measured, due to the degeneracy of the spin parameter $a$ and the tidal charge $b$ in Eq. (9). Indeed, the degeneracy in the leading-order term of the BZ power has been confirmed in a general stationary and axisymmetric BH spacetime Konoplya:2021qll.
What is really of concern, as suggested in Ref. Camilloni:2023wyn, is the fractional deviation of the high-spin factor $f_{\beta}$ with respect to its GR limit $f_{0}$, or equivalently, is the relative difference in BZ power between a Kerr and a braneworld BH with the same mass $M$ and angular velocity $\Omega_{\text{H}}$,
@@
It is clear from Fig. 6 that a positive $\beta$ always leads to a negative $\Delta\dot{E}$, and therefore tends to reduce the energy extraction rate. With a high-spin factor accurate to the second order in $\Omega_{\text{H}}$, the effect of positive tidal charges can suppress the BZ power of a braneworld BH by up to approximately $15.2\%$ compared to that of a Kerr BH. Conversely, a negative tidal charge acts to boost the power by $66.5\%$ at most from a purely theoretical perspective. Such an enhancement, however, can only be achieved at $\beta\approx-58.8$, a value that is likely not allowed in realistic astrophysical contexts. For example, observations of the BH shadow constrain the dimensionless tidal charge to $\beta=-0.1^{+0.6}_{-0.5}$ Banerjee:2019nnj; Banerjee:2022jog; Krishnendu:2024jkj, corresponding to a maximum $\Delta\dot{E}$ of $11.8\%$.
Figure 6: Percentage-scaled $\Delta\dot{E}$ as a function of the dimensionless angular velocity $\omega_{\text{H}}\equiv M\Omega_{\text{H}}$ for different values of $\beta$, where the independent variable is bounded by $\sqrt{1-\beta}/(2-\beta)$ according to Eq. (9). Since there are no allowed configurations outside the dashed gray line, $\Delta\dot{E}$ reaches its minimum of approximately $15.2\%$ at $\omega_{\text{H}}=0.392$ and $\beta=0.765$. The inset shows the envelope at extremely negative tidal charges, revealing a maximum enhancement $\Delta\dot{E}_{\text{max}}\approx 66.5\%$ for $\omega_{\text{H}}=0.127$ with $\beta=-58.818$. Configurations in the light green area are compatible with the BH shadow observations Banerjee:2019nnj; Banerjee:2022jog; Krishnendu:2024jkj
@@
which behaves qualitatively similarly to Eq. (76), as presented in Fig. 7. Moreover, the light green areas in Figs. 6 and 7 correspond to the allowed regions from the BH shadow observations for the relative energy flux $\Delta\dot{E}$ and the relative angular momentum flux $\Delta\dot{L}$, respectively.
Figure 7: Relative difference in angular momentum extraction rate, $\Delta\dot{L}$, between Kerr and braneworld Kerr BHs with identical $M$ and $\Omega_{\text{H}}$. The function (77) exhibits a maximum of $33.7\%$ at $\omega_{\text{H}}=0.143$, $\beta=-45.602$, and a minimum of $8.3\%$ at $\omega_{\text{H}}=0.401$, $\beta=0.749$. The light green area delineates the observational constraints $\beta\in[-0.6,0.5]$ Banerjee:2019nnj; Banerjee:2022jog; Krishnendu:2024jkj
@@ V Conclusions
For Kerr and braneworld BHs sharing the same mass $M$ and angular velocity $\Omega_{\text{H}}$, it is found that the BZ power is suppressed by the positive tidal charge, with a maximum reduction of approximately $15.2\%$ given our current level of precision, whereas a negative tidal charge can enhance the power by $66.5\%$ at most theoretically. A similar trend also holds for the angular momentum extraction rate, which is suppressed by a positive tidal charge and enhanced by a negative one, with the extremal values of $8.3\%$ and $33.7\%$, respectively. These results are helpful in placing independent constraints on the tidal charge parameter and the extra dimension size through future horizon-scale observations.